OpenStudy (anonymous):

Type the standard form of the equation of the circle with the center(-6,8),radius=10

OpenStudy (dumbcow):

\[(x-h)^2 + (y-k)^2 = r^2\] center = (h,k)

OpenStudy (anonymous):

Thank you!

OpenStudy (dumbcow):

yw :)